* 3. Steffensens Method is a combination of: (1 Point) Bisection method and Secant method Newtons method and the Aitkens A
Answer #1

Fixed point iteration and the Aitken's ◇2 method.

Suppose we have a fixed point iteration

P0, p1=g(p0), p2=g(p1)....

Once we have p0, p1, and p2, we can compute

P0`= p0- ((p1-p0)^2/(p2- 2p1+p0))

At this point we restart the fixed point iteration with p0=p0`

P3=p0`, p4=g(p3), p5=g(p4)

And compute p3` =p3 - ((p4-p3)^2)/(p5-2p4+p3)